The market demand curve p=30-Q there are only two firms in the market each with marginal cost of $0 what is market price in the stackelberg model


1
Expert's answer
2022-03-30T14:13:41-0400

TR= P.Q

= (30- Q)Q

= 30Q- Q"^2"

MR= 30- 2Q

At Equilibrium,

MR= MC

30- 2Q= 0

Q= 15

P= 30- 15

= 15